OpenStudy (anonymous):

solve for x -6(2x-10)+12x=180

OpenStudy (anonymous):

\[ -6(2x-10)+12x=180\] distribute the \(-6\) first what do you get?

OpenStudy (anonymous):

@satellite73 isnt the -6(2x-10) equals -12x or am i wrong

OpenStudy (anonymous):

you are wrong you need to do two multiplications

OpenStudy (anonymous):

ooh i get it what grade level problem is this??

OpenStudy (anonymous):

i got that its -12x-60+12x=180

OpenStudy (anonymous):

i never learned this.

OpenStudy (anonymous):

k good now combine like terms

OpenStudy (anonymous):

algebra 2 honors

OpenStudy (anonymous):

Oh ok i only take 9th grade algebra 1 honors

OpenStudy (anonymous):

then its... 0x=120 then what?

OpenStudy (anonymous):

actually you get \(-60=180\) because \(12x-12x=0\)

OpenStudy (anonymous):

Its not 0x=120 its just x = 120 am i wrong satellite?

OpenStudy (anonymous):

no \(0x=0\)

OpenStudy (anonymous):

so its just 120? or x=120

OpenStudy (anonymous):

but its -12x+12x which is 0x

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and \(0x=0\) right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

yeah so then that means its 0x=120

OpenStudy (anonymous):

no, there is still a \(-60\) on the left

OpenStudy (anonymous):

or 0=120

OpenStudy (anonymous):

Oh right

OpenStudy (anonymous):

actually \(-60=180\)

OpenStudy (anonymous):

which of course is wrong, so there is no solution to this equation

OpenStudy (anonymous):

does she have to simplify that or thats the official answer

OpenStudy (anonymous):

oh okay so the answer is no solution? Thanks so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

let sattelite get the medal he deserves it.